• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name adaq

    Log entry time 08:05:30 on October 22, 1998

    Entry number 3245

    keyword=Beam Time Accounting
    Hall A

    10/22/1998 0000-0800
    Experiment #: E94-010
    Shift Leader: Deur A.
    Shift Workers: Choi S.

    00000100020003000400050006000700Total Time
    Acceptable Beam
    in Use
    0 0 0 0 0 0 0 0 0
    Beam Available
    but Not in Use
    0 0 0 0 0 0 0 0 0
    Beam Not Available
    or Unacceptable
    0 0 0 0 0 0 0 0 0
    Accelerator Con-
    figuration Chage
    1 1 1 1 1 1 1 1 8
    Experiment
    Ready
    0 0 0 0 0 0 0 0 0
    Planned Config-
    uration Change
    1 1 1 1 1 1 1 1 8
    Unplanned
    Experiment Down
    0 0 0 0 0 0 0 00
    Hall not
    ready reasons
    Target 0 0 0 0 0 0 0 0 0
    HRSe 0 0 0 0 0 0 0 0 0
    HRSh 0 0 0 0 0 0 0 0 0
    Data Acq 0 0 0 0 0 0 0 0 0
    Other 0 0 0 0 0 0 0 0 0

    Comments: 00:25 NMR measurement. Target Pol. 31.2-43 % 2:00 Begin to go in the Compton Chicane 2:30 NMR: 32.9-44.7 % 3:15 hadron VDCs trip 5:40 End of the Compton test. NMR: 34.5-48.6 6:00Controled access 7:NMR: 35.11-47.7 %

    Data Entry by: Deur A.